OpenStudy (anonymous):

Sean, a freelance editor, charges the rates shown in the table below to edit manuscripts. The cost per page increases as the quality of editing improves. Sean also gives a 5% discount if the entire amount is paid up front. Type Of Editing Cost Per Page Express Proofreading $2.00 Basic Proofreading $2.95 Extended Proofreading $3.95 Deep Editing $11.00 Michelle has an 85-page manuscript. The equation below shows the relationship between the total cost of editing 85 pages, T, and the cost per page, c, if she gets the 5% discount.

OpenStudy (anonymous):

Michelle has an 85-page manuscript. The equation below shows the relationship between the total cost of editing 85 pages, T, and the cost per page, c, if she gets the 5% discount. 85c - 0.05(85c) = T Using the equation, what is the best quality of editing that Michelle can get done for a maximum of $161.50? A. Express proofreading B. Basic proofreading C. Extended proofreading D. Deep editing

OpenStudy (jacob902):

You are trying to find c if you know T. Given: 85c - 0.05(85c) = T (1 - 0.05) x 85c = T 0.95 x 85c = T 80.75c = T 80.75c = 161.50 c = 161.50 / 80.75 = 2.00 She can get $2.00 worth of editing per page.

OpenStudy (anonymous):

so A

OpenStudy (jacob902):

yes

OpenStudy (anonymous):

thnx :)